OpenStudy (anonymous):

Ten grams of a gas occupies 12.5 liters at a pressure of 42.0 cm Hg. What is the volume when the pressure has increased to 75.0 cm Hg? 6.00 L 7.00 L 8.00 L none of the above

OpenStudy (anonymous):

Since the question doesn't say anything about the mass changing we use the equation: \[P _{1}V _{1} = P _{2}V _{2}\] 12.5*42 = 75*P2 P2 = 7.00 L

OpenStudy (anonymous):

thankx....